OpenStudy (erinweeks):

In , is a right angle. Find the remaining sides and angles. Round your answers to the nearest tenth. a = 3, c = 19 a. = 9.1°, = 80.9°, b = 18.8 c. = 14.5°, = 75.5°, b = 18.8 b. = 80.9°, = 9.1°, b = 18.8 d. = 75.5°, = 14.5°, b = 18.8

OpenStudy (phi):

a picture would help

OpenStudy (anonymous):

Pythagorean theorem. c^2 -a^2=b^2

OpenStudy (anonymous):

\[19^2-3^2=c^2\]

OpenStudy (anonymous):

361-9=352

OpenStudy (anonymous):

\[352=c^2\]

OpenStudy (anonymous):

\[\sqrt{352}\]=c

OpenStudy (anonymous):

c=18.76166304

OpenStudy (anonymous):

However, this is not in the correct form. it gives you 18.8=c

OpenStudy (anonymous):